y′′+ 4y= sin 5x+x−1

1
Expert's answer
2022-03-02T16:58:24-0500

The characteristic equation "k^2+4=0" has roots "k_1=2i" and "k_2=-2i." Therefore, the general solution of the differential equation is of the form:

"y=C_1\cos 2x+C_2\sin 2x+y_p,"

where "y_p=a\cos 5x+b\sin 5x+cx+d." It follows that

"y_p'=-5a\sin 5x+5b\cos 5x+c,
y_p''=-25a\cos 5x-25b\sin 5x"

Then

"-25a\cos 5x-25b\sin 5x+4(a\cos 5x+b\sin 5x+cx+d)=\sin 5x +x-1."

We get the following equation

"-21a\cos 5x-21b\sin 5x+4cx+4d=\sin 5x +x-1,"

and hence

"-21a=0,\ -21b=1,\ 4c=1,\ 4d=-1."

It follows that "a=0,\ b=-\frac{1}{21},\ c=\frac{1}4,\ d=-\frac{1}4."

We conclude that the general solution of the differential equation is of the form

"y=C_1\cos 2x+C_2\sin 2x-\frac{1}{21}\sin 5x+\frac{1}4x-\frac{1}4."